The taxi charges in a city consist of a fixed charge together with the charge for the distance covered. For a distance of 10 km, the charge paid is Rs 105 and for a journey of 15 km, the charge paid is Rs 155. What are the fixed charges and the charge per km? How much does a person have to pay for travelling a distance of 25 km?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Here is the solution.
Let the fixed charge be Rs x and per km charge be Rs y
According to the question,
x + 10y = 105 …………….. (1)
x + 15y = 155 …………….. (2)
From (1)
we get
x = 105 – 10y ………………. (3)
Substituting the value of x in (2),
we get
105 – 10y + 15y = 155
5y = 50
y = 10 …………….. (4)
Putting the value of y in (3)
we get
x = 105 – 10 × 10 = 5
Hence, fixed charge is Rs 5 and per km charge = Rs 10
Charge for 25 km = x + 25y = 5 + 250 = Rs 255

Thanks